After 12 years disha shall be 3 times as oldas 4 years ago . find her present age

Grade:8

1 Answers

Arun
5 years ago
Ket present age of disha is x years. Hence...(x+12) = 3 (x-4) ......x+12 = 3 x-12....2x = 24....x = 12 years